OpenStudy (lilcathyx):

i just want to know if i did it correct... thats all ! :) 1. The main tank has a radius of 70 feet. What is the volume of the quarter-sphered sized tank? Round your answer to the nearest whole number

OpenStudy (lilcathyx):

V=4/3 x 3.14 x r^3 V=1.3 x 3.14 x 70^3 V= 1.3 x 3.14 x 343000 V= 1,400,126 Is that already in cubic feet or no ?

OpenStudy (nothingwasthesame):

um i would say yes but to figure it out we must start at the beginning

OpenStudy (lilcathyx):

okkk :o

OpenStudy (nothingwasthesame):

lol what are you trying to figure out in all and all

OpenStudy (lilcathyx):

if my final answer is in cubic feet ? because f its not idk how to put it in cubic feet..

OpenStudy (nothingwasthesame):

oh yes

OpenStudy (lilcathyx):

it is in cubic feet?

OpenStudy (nothingwasthesame):

no it is not because i do not see where you put in cubic feet unless the work already was i cubic feet

OpenStudy (lilcathyx):

i dont understand that much that whole cubic feet thingy. thats why im asking if its already in cubic feet or if there's another step to do to make it in cubic feet//

OpenStudy (nothingwasthesame):

ok will first go to this website and read http://www.thecalculatorsite.com/articles/units/calculating-cubic-conversions.php tell me if it helps then if it does not well start from scratch k :)

OpenStudy (lilcathyx):

okk brb :)

OpenStudy (lilcathyx):

im sorry but im just plain lost now.

OpenStudy (nothingwasthesame):

lol ok well first explain the question and the problem

OpenStudy (lilcathyx):

i have to find the volume of a sphere. the volume of a sphere is V=4/3 x 3.14 x r^3

OpenStudy (lilcathyx):

@jim_thompson5910 help pleasee ?

jimthompson5910 (jim_thompson5910):

first thing, I would keep 4/3 as it is and not use 1.3 as that is a really bad approximation. If you need to use an approximation, use more decimal places (so maybe 4/3 = 1.333) but if I were doing this problem, I would simply type (4/3)*pi*70^3 into a graphing/scientific calculator as one full expression. That way there isn't as much rounding error

jimthompson5910 (jim_thompson5910):

also, they want the volume of the "quarter-sphered sized tank" so you'll find the volume of the entire sphere and then divide that result by 4

OpenStudy (lilcathyx):

it gave me this --> 1436026.6666666665

jimthompson5910 (jim_thompson5910):

now divide that by 4

OpenStudy (lilcathyx):

divided by 4 is equal to this -->359006.6666666666 rounding to the hundredth would be 359007 right ?

jimthompson5910 (jim_thompson5910):

"Round your answer to the nearest whole number" so you'll go from 359006.6666666666 to 359007 you are correct

jimthompson5910 (jim_thompson5910):

the volume units are in cubic feet

OpenStudy (lilcathyx):

the final answer will be 359007 cubic feet ? :O

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (lilcathyx):

Ahhhhhhhhh ! Thankkk youuuu ! ^.^

jimthompson5910 (jim_thompson5910):

you're welcome